OpenStudy (anonymous):

Suppose that (x, (-24/25)) is a point in quadrant IV lying on the unit circle. Find x . Write the exact value, not a decimal approximation. what does this mean? please help me

OpenStudy (luigi0210):

I think it means we need to solve.. So on the unit circle the points are: (cosx, sinx) right? So far we have sinx. which is opp/hyp Now we just need cosx which is adj/hyp:  So find x, using a^2+b^2+c^2

OpenStudy (campbell_st):

well just draw a picture 

OpenStudy (luigi0210):

The x would be better as b..

OpenStudy (anonymous):

so -24^2 +x^2 = 25^2?

OpenStudy (luigi0210):

Yes, and that will give us the adj angle we need

OpenStudy (anonymous):

so x =7?

OpenStudy (luigi0210):

Yup, so it would be cosx=7/25.. cos/sec are positive in the 4th and 1st quad

OpenStudy (anonymous):

so since it says to just find x, that means I should say x = 7 right?

OpenStudy (campbell_st):

no, the circle has a radius of 1 unit... hence the name, unit circle the the value of x is 7/25 hope that makes sense...

OpenStudy (luigi0210):

Uhm, yea.. what campbell said xD

OpenStudy (luigi0210):

Fast typer xD

OpenStudy (anonymous):

k thanks you two:D